OpenStudy (anonymous):

please help! i need to solve for x in the problem: 3^(2x)-2*3^(x+5) + 3^10 =0

OpenStudy (anonymous):

\[3^{2x}-2*3^{x+5} + 3^{10} = 0\] Is it right?

OpenStudy (anonymous):

yes, hope so

OpenStudy (anonymous):

yeh thats how its written

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

3^x=t t^2-2*3^5t+3^10=0

OpenStudy (anonymous):

So then do I combine the "-2*3^5t ?"

OpenStudy (anonymous):

just complete the square and you'll get (t-3^5)^2=0

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

i think i need help with the steps?

OpenStudy (anonymous):

please?

OpenStudy (anonymous):

Like uzma said use a substitution t=3^x. First notice that 3^(x+5)=(3^x)*(3^5)=(3^5)*y. Then the equation reduces to t^2-2(3^5)y+3^10=0. Then solve this as a quadratic in y, using 'complete the square'. Make sense?

OpenStudy (anonymous):

I used y so when you see y think of it as t

OpenStudy (anonymous):

before i complete the square do i simplify to t^2-6^5t+3^10?

OpenStudy (anonymous):

oh i get it thank you